Identify the level curves of the following functions:

(i) √(x2+y2)

(ii) √(4 - x2 + y2)

(iii) x-y

(iv) x/y


1
Expert's answer
2022-02-01T02:06:28-0500

A level curve of a function "f(x,y)" is the curve with the equation "f(x,y)=c", where "c" is an arbitrary constant.

(i) The level curves of the function "\sqrt{x^2+y^2}" can be determined by the equations:

"\sqrt{x^2+y^2}=r,\, r\geq0"

"x^2+y^2=r^2" - this is the equation of the circle centered at (0,0) and with radius "r".

Therefore, the level curves of the function "\sqrt{x^2+y^2}" is a family of circles centered at (0,0).


(ii) The level curves of the function "\sqrt{4 - x^2 + y^2}" can be defined by the equations:

"\sqrt{4 - x^2 + y^2}=c,\, c\geq 0"

"- x^2 + y^2=c^2-4"

"(y+x)(y-x)=c^2-4"

We obtained the family of hyperbolas with asymptotes "y-x=0" and "y+x=0" (corresponding to "c\ne2"), and the curve "y^2-x^2=0" (corresponding to "c=2"), which is a union of two lines "y-x=0" and "y+x=0"


(iii) The level curves of the function "x-y" can be defined by the equations:

"x-y=c" . These are all straight lines which are parallel to the line "x-y=0".


(iv) The level curves of the function "x\/y" can be defined by the equations:

"x\/y=c", or "x=cy,\,y\ne0" - this is a family of open rays converging at point (0,0), excluding two rays with "y=0".